# MixLM: High-Throughput and Effective LLM Ranking via Text-Embedding Mix-Interaction

### v1

@@ 5.2.1. Text-Embedding Mixed Input
Common open-source LLM inference engines focus on pure-text inputs, restricting requests to tokenized sequences. To enable mix-interaction in MixLM, we extended the input interface to additionally accept precomputed embedding tensors. These embeddings are transmitted using a structured feature payload abstraction added to the gRPC request schema. Each feature payload encapsulates a binary buffer containing tensor values, along with metadata specifying the tensor’s shape, element type, and optional serialization format (e.g., NumPy (numpy)-compatible). This design allows clients to include dense embeddings alongside text inputs without modifying the text-based interface or introducing format-specific logic in the serving stack.
@@
On the server side, the preprocessing module decodes and validates these feature payloads into native tensor objects, applying metadata-driven reshaping and casting as necessary. The resulting tensors are then attached to the model input context for downstream inference. This approach enables direct reuse of the embeddings computed by upstream retrieval or feature pipelines, avoiding redundant re-computation and reducing latency. The extended schema remains fully backward-compatible with existing text-only requests while introducing minimal overhead in serialization and deserialization.
@@
Table 2. Attention and linear-layer compute under different conditions, in terms of $(T_{q},T_{i},N_{i})$, where $T_{i}$ denotes the full item-text length and $K$ is the MixLM compression factor (item length becomes $T_{i}/K$ in the prompt). The symbol $\propto$ indicates proportionality up to architecture-dependent constants (layers, heads, hidden size, etc.). The bottom blocks instantiate the expressions for $N_{i}=250$ and $K=450$.
@@
Amortized prefill ++ MixLM (item length $T_{i}/K$) $\propto T_{q}^{2}+N_{i}\bigl(2(T_{i}/K)T_{q}+(T_{i}/K)^{2}\bigr)$ $\propto T_{q}+N_{i}T_{i}/K$
Naive (full-text items) with $N_{i}=250$ $\propto 250(T_{q}+T_{i})^{2}$ $\propto 250(T_{q}+T_{i})$
Amortized prefill (full-text items) with $N_{i}=250$ $\propto T_{q}^{2}+250(2T_{i}T_{q}+T_{i}^{2})$ $T_{q}^{2}+500T_{i}T_{q}+250T_{i}^{2}$ $\propto T_{q}+250T_{i}$
Amortized prefill ++ MixLM with $N_{i}=250$, $K=450$ $\propto T_{q}^{2}+250\bigl(2(T_{i}/450)T_{q}+(T_{i}/450)^{2}\bigr)$ $T_{q}^{2}+\frac{10}{9}T_{i}T_{q}+\frac{1}{810}T_{i}^{2}$ $\propto T_{q}+250T_{i}/450$ $T_{q}+\frac{5}{9}T_{i}$
5.2.2. Shared-Prefix Prefill Optimization
In large-scale search and recommendation platforms, a single ranking request typically contains hundreds to thousands of candidate items, all associated with the same query and user-history features. After we compress each item’s description into a small, fixed set of embedding tokens, the computational burden of inference shifts predominantly to the repeating query/user-history side, because these tokens now constitute the majority of the prompt. For example, in our production system a candidate item was previously represented by roughly 900 tokens as median; with MixLM this representation is reduced to only two tokens—one special token and one embedding token.
To improve our GPU utilization during online inference, we use a batched inference mechanism. That is, for each query, we batch several prompts, each corresponding to a different candidate item, and run inference on the batch. Given that all prompts in a batch share the same prefix, this allows for reusing the KV cache across different prompts. This is particularly important when combined with the mixed-input system: once the item representation is reduced to just a few tokens, the shared query-prefix computation dominates the overall inference cost.
Let $T_{q}$ and $T_{i}$ denote the number of shared query tokens and the number of item tokens in the full-text representation, respectively, and let $N_{i}$ denote the ranking depth. Each transformer block consists of a multi-head self-attention module and a set of linear layers (query/key/value projections and a feed-forward network). For a sequence of length $L$, we write the attention and linear costs using proportionality (ignoring architecture-dependent constants such as depth, number of heads, and hidden/FFN dimensions) as
$\text{Self-attention cost}\;\propto\;L^{2},\qquad\text{Linear/FFN cost}\;\propto\;L.$
Thus, attention is quadratic in the sequence length $L$, while the linear layers are linear in $L$ but may have a larger constant factor.
Naïve prefill cost (full-text items).
Without amortization, each candidate item forms an independent sequence of length $(T_{q}+T_{i})$. Across $N_{i}$ items, the total attention and linear prefill costs scale as
@@ 5.2.2. Shared-Prefix Prefill Optimization
When both $T_{q}$ and $T_{i}$ are large, $F_{\text{att}}^{\text{naive}}$ contains dominant contributions from $N_{i}T_{q}^{2}$, $N_{i}T_{i}^{2}$, and $N_{i}T_{q}T_{i}$, while $F_{\text{lin}}^{\text{naive}}$ grows linearly in $(T_{q}+T_{i})$.
Amortized prefill (full-text items).
With amortized prefill, only the first sequence in a batch performs the full prefill over the query prefix; the remaining $N_{i}-1$ sequences reuse its KV cache and do not recompute query self-attention. The attention cost decomposes into
[…]
while the linear layers are applied once to the query prefix and once to each item token:
$F_{\text{lin}}^{\text{amort, full}}\;\propto\;T_{q}+N_{i}T_{i}.$
Amortized prefill with MixLM compression.
In MixLM, we use $T_{i}$ to denote the length of the original full item text, and introduce a compression factor $K>1$ such that the item is represented by $T_{i}/K$ tokens in the mixed-input prompt. In our online A/B test, we obtained a compression factor of $K\approx 450$, i.e., each item is compressed from hundreds of tokens down to two tokens in the prompt.
Under amortized prefill with MixLM, the query prefix of length $T_{q}$ is still processed once, but each item now contributes only $T_{i}/K$ tokens. The attention cost becomes
@@
Compared to the full-text setting, both the attention and linear components of the item-dependent cost are reduced by approximately a factor of $K$ on the item side. In particular, the dominant cross-term in attention changes from $N_{i}T_{i}T_{q}$ to $N_{i}(T_{i}/K)T_{q}$, and the item-side linear cost changes from $N_{i}T_{i}$ to $N_{i}(T_{i}/K)$.
To realize this amortization, we explored two approaches:
@@
  • In-batch prefix caching We reuse the key–value (KV) states from the first prompt so that all items in the batch share the same query-prefix computation in a single forward pass. The attention computation merges two operations:
    1. (1)
      suffix tokens attend to all prefix tokens from the first prompt via dense paged attention;
    2. (2)
      suffix tokens attend to themselves via regular causal attention.
  • Multi-item scoring We score multiple items against a single query within one forward pass by concatenating the query and all item texts into a single sequence (sglang_pr_10979), separated by a special delimiter token. FlashInfer’s item-aware masking ensures that each item attends only to the query and not to any other item. During prefill, metadata such as prefix length, item boundaries, and maximum item length is passed directly to the FlashInfer kernel to build the correct attention mask and prevent cross-item dependencies.
Because $T_{i}$ is extremely small in our mixed-input approach, amortized prefill dramatically reduces redundant prefix computation and yields substantial inference-throughput improvements in production.
5.2.3. CPU Overhead Reduction
Multi-Process gRPC
@@ 5.2.3. CPU Overhead Reduction
After optimizing both the model and the SGLang engine, we observed that request preprocessing in the Python gRPC service became the new performance bottleneck. Although the gRPC server is implemented with asynchronous Python, tasks such as request deserialization, feature decoding, and context construction remain constrained by Python’s Global Interpreter Lock (GIL) and the single-threaded nature of the asyncio event loop. High request volume can exacerbate these bottlenecks due to contention on the event loop and per-request overhead, limiting throughput despite asynchronous I/O. To overcome this constraint, we implemented a multiprocessing server architecture that decouples the gRPC server from the SGLang engine. This design enables preprocessing and request handling to run across multiple CPU cores in parallel, while the SGLang engine executes independently. As a result, this architecture yields a 40% improvement in end-to-end throughput.
Figure 4. Multiprocessing server for gRPC
Batch Send.
@@
To better leverage amortized prefill, the SGLang scheduler must group prompts that share the same query. To address this, we introduced an explicit batch send pathway in which the entire set of item prompts for a query is serialized and sent as a single ZMQ message. This guarantees that the scheduler receives and processes the full batch together, allowing it to admit the batch as a unit and apply in-batch prefix caching efficiently.
Multi-Process Pallellization
@@
Given our workload characteristics, the prefill phase completes rapidly on GPU, but request scheduling on a single CPU core became the bottleneck. We addressed this by deploying multiple serving engine’s scheduler processes (e.g., 5 workers) sharing a single GPU instance, with each process allocated a dedicated memory partition (e.g., 20% GPU memory per worker). This parallelizes CPU scheduling overhead across multiple cores while maintaining full GPU utilization, effectively pipelining batch preparation with GPU execution.
@@ 5.2.4. Eliminating Redundant KV-Cache Operations
SGLang’s default behavior caches computed key-value pairs during prefill to enable efficient autoregressive decoding. However, our scoring use case is prefill-only with no decode phase. We modified the cache management to immediately free KV cache memory upon request completion rather than persisting it for non-existent decode operations. This eliminates unnecessary memory allocation, cache eviction logic, and pool management overhead on the critical path. More importantly, it increases effective GPU memory capacity, allowing higher batch sizes and concurrent request throughput.
@@ 6. Experimental Results
We perform offline and online experiments to understand the ranking quality and throughput improvements of our proposed framework. We use a held-out test set, that has been labeled using our internal LLM judge.
We compare our mixed input system with a few baselines. The first baseline is a bi-encoder embedding-based retrieval model, which has been distilled from our LLM relevance judge. This is the retrieval layer of our semantic search stack. The next baseline is a pure-text system, that is, the 0.6B model available from our second training stage. This represents the ideal ranking model, though it is too inefficient to be deployed. In addition, we compare with LinkedIn’s limited-scale production baseline (behdin2025scaling) that uses a near-line text-summarization model to summarize item descriptions and apply aggressive pruning to the ranker model.
In terms of metrics, we report the NDCG@10 on the labeled held-out set. Additionally, we report the maximum throughput we can achieve for each baseline in our production setup. In particular, we report the maximum number of items we can score per GPU per second, while ensuring the P99 end-to-end latency remains below 500 milliseconds.
6.1. Comparisons with Baselines
Table 3. Rankin quality comparison of several ranking approaches: We report NDCG@10 with respect to the labels provided by our relevance judge. See Section 6 for a description of baselines.
@@ 6.1. Comparisons with Baselines
Table 4. GPU efficiency comparison of ranking approaches. For each ranker, we report the observed maximum throughput (items scored per H100 GPU per second) under a fixed 500 ms latency budget. For reference, we also include the throughput of the retrieval layer (embedding-based exhaustive GPU retrieval, deployed on A100 GPUs). See Section 6 for details on the baseline methods.
Model QPS(Items/Sec/GPU) Latency (ms) GPU
Full Text 290 ¡500 H100
Summarized, Pruned 2 200 ¡500 H100
MixLM 22 000 ¡500 H100
Embedding Retrieval $1.6\times 10^{9}$ ¡100 A100
Table 3 summarizes the ranking quality and Table 4 summarizes system efficiency of the baselines we discussed. Embedding-based retrieval, while able to process well over $1.6\times 10^{9}$ documents per seconds on a single weaker A100 GPU, delivers the weakest ranking quality and is therefore usually employed only as a coarse first-stage filter. Feeding the full documents to a large language model (pure-text LLM) pushes NDCG to 0.9432, but its heavy token footprint limits throughput to merely 290 items/sec/GPU. Compressing each document into a short summary before prompting the LLM (summarized text LLM) and pruning ranker LLM’s parameters by  37% (behdin2025scaling) alleviates the latency bottleneck somewhat—throughput rises to  2,200 items/sec/GPU—yet ranking quality NDCG@10 is still an acceptable value as 0.9218. Our proposed MixLM achieves a similar NDCG@10 of 0.9239. However, we observe that MixLM allows us to scale up the throughput to 22,000 items/sec/GPU, i.e., a 10.0x speed-up over summarized-text LLM and a 75.9x speed-up over full-text LLM, at the cost of only a 1.8-point NDCG gap with respect to the latter. In other words, MixLM hits a sweet spot in the quality–efficiency trade-off: it is orders of magnitude faster than existing LLM-ranking baselines while retaining virtually the same retrieval effectiveness, making it an attractive drop-in replacement for latency-critical production systems.
@@ 6.2. Online A/B Test
Table 5 summarizes the findings. After rolling out MixLM in LinkedIn’s AI Job Search and running extensive A/B tests, we observed: MixLM is has on-par metrics comparing with the existing limited-scale production baseline(summarized text LLM)  (behdin2025scaling). With the 10.0x gain in serving throughput under the same latency budget, MixLM empowered the full-traffic rollout of LLM-ranking-based Semantic Job Search and for the first time, consequently producing a significant 0.47% increase in Daily Active Users, compared with the traditional job search.
Table 5. Online A/B Test
Experiment Group DAU
@@
Semantic Job Search (powered by MixLM) +0.47%
@@ 6.3. Ablation Analysis
We conducted extensive ablation studies on a smaller subset that mirrors the distribution of the full training corpus, which is common practice in LLM training because full-scale experiments would have required high GPU hours. Within this setting, we systematically analyzed every major component of MixLM training—(1) dataset size, (2) number of token embeddings per item, (3) domain-reasoning fine-tuning, (4) auxiliary loss functions, and (5) curriculum learning strategies. Also, we conducted ablations on different inference engine optimization methods aimed at improving the computation of shared prefix prompt (query and the user-history).
@@ 6.3.1. Scaling Up
Expanding the amount of training data had a clear positive effect on ranking quality. In the ablation study, we progressively enlarged the corpus from 160 K to 1.08 M instances, observing a consistent increase in NDCG@10 at each step. Motivated by this trend, the model deployed in the online A/B experiment was trained on the full dataset of 10.9 M samples.
A similar monotonic gain was observed when we varied the representational granularity of each item. Raising the number of embedding tokens per item from 1 to 50 produced steady improvements in NDCG@10. Nonetheless, to meet strict latency requirements, the production configuration retains a single embedding token per item (followed by 1 special end-of-sentence text token).
Table 6. Ablations on Dataset Size, on small sub-dataset
Num Training Sample NDCG@10
@@
Table 7. Embeddings Token Number per Item, on small sub-dataset
Tokens/item NDCG@10
@@ 6.3.2. Domain Reasoning Fine-Tuning
The ranker LLM serves as the backbone of our model architecture: both the query tokens and the compressed embedding tokens pass through this ranker to produce the final binary (yes/no) prediction. Consequently, the quality of the ranker LLM directly determines the overall ranking performance.
We found that employing a stronger base model for the ranker substantially improves the final results. Specifically, we compared two configurations: (1) using a open-source pretrained 0.6B model as the base, and (2) using a domain reasoning fine-tuned version of the same model. The results, summarized in Table 8, clearly show that the domain reasoning tuned base model yields superior bi-encoder performance.
Table 8. Effect of Domain Reasoning Fine-tuning on ranker LLM’s Ranking performance, on small sub-dataset
Ranker Base Model NDCG@10
vanilla open-source LLM
@@
Based on these findings, we have adopted Domain Reasoning Fine-Tuning as the first stage of our training recipe. In the first stage, we obtain a more powerful base model for the ranker LLM via task-specific COT data, guided by our internal 7B relevance judge LLM as the teacher model. After that, we use this tuned model as the Ranker LLM’s starting base model, to initialize the training of stage III: Joint Ranking Training
@@ 6.3.3. Self-Alignment and Auxiliary Losses
To better understand the contribution of each auxiliary loss, we conducted an ablation study by selectively disabling different components. In all cases, we used the same setup (small sub-dataset, 50 embedding tokens per item) and varied the presence of distillation loss and self-alignment loss.
Table 9. Ablation study on auxiliary losses: 50 embedding tokens per item, on small sub-dataset
Setup NDCG@10
@@
+self-alignment loss +0.0014
+distillation loss +0.0091
+self-alignment +distillation losses +0.0108
We observe that:
  • Distillation loss contributes a significant improvement (+0.009 relative to KL-only).
  • Self-alignment alone provides only marginal gains over the KL-only baseline.
  • Combining distillation and self-alignment achieves the best performance (NDCG@10 +0.0108), confirming that the two auxiliary losses are complementary.
These results highlight the importance of including distillation in addition to KL and self-alignment losses to maximize ranking quality.
6.3.4. Curriculum Learning Strategy
In addition to the auxiliary loss components, we investigated the effect of different phased training strategies that align with curriculum learning principles for Stage III in Section 4.3.3. Curriculum learning suggests that training with carefully sequenced objectives or loss weightings can lead to better convergence and final model performance. In the context of text-embedding alignment problem, we explored the following curriculum strategies:
  • No Curriculum Learning (Baseline): focus more on the KL divergence loss but less on distillation loss and self-alignment regularization.
  • Two-Phase Curriculum Strategy: On top of the baseline, add an alignment phase at the beginning that has high self-alignment weights with low KL divergence weights.
  • Three-Phase Curriculum Strategy: Add an intermediate phase to two-phase curriculum, that has equal weighting across all three loss objectives (distillation, SFT, and self-alignment).
As seen in Table 10, curriculum learning improves ranking performance, indicating the alignment is necessary. The two-phase curriculum outperforms the three-phase approach, suggesting that a direct transition between alignment-focused and task-performance-focused training is more effective than including an intermediate balanced phase. This finding indicates that once the ranker-encoder alignment is sufficiently established, the model can more directly optimize for the downstream ranking task without the gradual intermediate phase.
Table 10. Ablation study on curriculum learning strategies, on small sub-dataset
Curriculum Strategy NDCG@10
no curriculum learning (Task)
two-phase (Alignment $\to$ Task) +0.0020
three-phase (Alignment $\to$ Balanced $\to$ Task) +0.0015

### v2

@@ 5.1. Text–Embedding Mixed Input
To support MixLM’s mixed-interaction design, we extend the serving interface to accept precomputed embedding tensors alongside text. Embeddings are transmitted via a structured feature payload added to the gRPC request schema, consisting of a binary tensor buffer and metadata specifying shape, data type, and optional serialization format (e.g., NumPy (Harris et al., 2020)).
@@
On the server side, the preprocessing module decodes and validates feature payloads into native tensors using metadata-driven reshaping and casting, and attaches them to the model input context. This enables direct reuse of embeddings produced by upstream pipelines, avoiding redundant computation and reducing latency. The schema remains fully backward-compatible with text-only requests and introduces minimal serialization overhead.
@@
Table 2. Attention and linear-layer compute under different conditions, in terms of $(T_{q},T_{i},N_{i})$, where $T_{i}$ denotes full item-text length and $K$ is the MixLM compression factor (item length becomes $T_{i}/K$). $\propto$ denotes proportionality up to architecture-dependent constants. Bottom blocks instantiate $N_{i}=250$, $K=450$.
@@
Amortized prefill ++ MixLM $\propto T_{q}^{2}+N_{i}\bigl(2(T_{i}/K)T_{q}+(T_{i}/K)^{2}\bigr)$ $\propto T_{q}+N_{i}T_{i}/K$
Naive, $N_{i}=250$ $\propto 250(T_{q}+T_{i})^{2}$ $\propto 250(T_{q}+T_{i})$
Amortized, $N_{i}=250$ $T_{q}^{2}+500T_{i}T_{q}+250T_{i}^{2}$ $\propto T_{q}+250T_{i}$
Amortized ++ MixLM $N_{i}=250$, $K=450$ $T_{q}^{2}+\frac{10}{9}T_{i}T_{q}+\frac{1}{810}T_{i}^{2}$ $T_{q}+\frac{5}{9}T_{i}$
5.2. Shared-Prefix Prefill Optimization
In large-scale ranking, each request typically scores hundreds to thousands of items sharing the same query and user context. After MixLM compresses each item into a small number of embedding tokens (from a median of $\sim$900 tokens to two tokens in production), inference cost becomes dominated by repeated query-prefix computation.
To improve GPU utilization, we batch candidate-item prompts per query and reuse the KV cache across prompts with a shared prefix. This is particularly effective under MixLM, where item tokens are minimal and query-prefix computation dominates cost.
Let $T_{q}$ be the query-prefix length, $T_{i}$ the full-text item length, and $N_{i}$ the ranking depth. Ignoring architecture-dependent constants, attention and linear costs scale as
$\text{Attention}\propto L^{2},\qquad\text{Linear}\propto L.$
Naïve prefill (full-text).
Each item is processed independently with length $(T_{q}+T_{i})$:
@@ 5.2. Shared-Prefix Prefill Optimization
Amortized prefill (full-text).
Query-prefix attention is computed once and reused:
$F_{\text{att}}^{\text{amort, full}}\propto T_{q}^{2}+N_{i}(2T_{i}T_{q}+T_{i}^{2}),\qquad F_{\text{lin}}^{\text{amort, full}}\propto T_{q}+N_{i}T_{i}.$
Amortized prefill with MixLM.
With item compression factor $K$, each item contributes $T_{i}/K$ tokens:
@@
Thus, item-side attention and linear costs are reduced by approximately $K$.
We implement amortization using two mechanisms:
@@
  • In-batch prefix caching: We reuse the key–value (KV) states from the first prompt so that all items in the batch share the same query-prefix computation in a single forward pass. The attention computation merges two operations: (1)suffix tokens attend to all prefix tokens from the first prompt via dense paged attention; (2)suffix tokens attend to themselves via regular causal attention.
  • Multi-item scoring: Multiple items are concatenated into a single sequence separated by delimiters (Sundara Raman Ramachandran, 2025). FlashInfer applies item-aware masking to prevent cross-item attention.
Because MixLM reduces $T_{i}$ to only a few tokens, shared-prefix amortization eliminates most redundant computation and delivers substantial inference-throughput gains in production.
5.3. Inference Engine CPU Overhead Reduction
5.3.1. Multi-Process gRPC
@@
After optimizing the model and the SGLang inference engine, request preprocessing in the Python gRPC service emerged as the dominant bottleneck. Although the gRPC server uses asynchronous Python, request deserialization, feature decoding, and context construction are constrained by the Global Interpreter Lock (GIL) and the single-threaded asyncio event loop. Under high request rates, event-loop contention and per-request overhead limit throughput despite asynchronous I/O.
To address this, we adopt a multiprocessing server architecture that decouples the gRPC frontend from the SGLang engine. Request handling and preprocessing are parallelized across multiple CPU processes, while the SGLang engine runs independently. This design fully utilizes multi-core CPUs and improves throughput by 40%.
5.3.2. Batch Send
@@
To fully exploit amortized prefill, the SGLang scheduler must admit prompts sharing the same query as a single batch. We therefore introduce an explicit batch send mechanism, where all item prompts for a query are serialized and transmitted as a single ZMQ message. This ensures that the scheduler processes the batch atomically and can reliably apply in-batch prefix caching.
5.3.3. Multi-Process Parallelization
@@
While GPU prefill executes efficiently, request scheduling on a single CPU core became a secondary bottleneck. We mitigate this by running multiple scheduler processes (e.g., five) per GPU, each assigned a dedicated GPU memory partition (e.g., 20% per worker). This parallelizes CPU-side scheduling while maintaining high GPU utilization, effectively pipelining batch preparation with GPU execution.
@@ 5.3.4. Eliminating Redundant KV-Cache Operations
SGLang’s default KV-cache behavior is designed for autoregressive decoding, which is not used in our prefill-only scoring workload. We modify cache management to immediately release KV cache memory after request completion, eliminating unnecessary allocation, eviction, and pool management. This both reduces CPU overhead on the critical path and increases effective GPU memory capacity, enabling larger batch sizes and higher concurrent throughput.
@@ 6. Experimental Results
We conduct both offline and online experiments to evaluate the ranking quality and serving efficiency of MixLM. Offline evaluation uses a held-out test set labeled by our internal LLM relevance judge. We compare against several baselines: (1) an embedding-based bi-encoder retrieval model distilled from the LLM judge (the first-stage retriever in our stack), (2) a pure-text 0.6B LLM ranker from the second training stage, representing an upper-bound model that is impractical to deploy, and (3) LinkedIn’s limited-scale production baseline that applies near-line text summarization and aggressive ranker pruning (Behdin et al., 2025).
We report NDCG@10 for ranking quality and maximum throughput measured as items scored per GPU per second, under a fixed p99 end-to-end latency constraint of 500 ms.
6.1. Efficiency and Relevance
Table 3. Ranking quality comparison of several ranking approaches. We report NDCG@10 with respect to labels provided by our relevance judge.
@@ 6.1. Efficiency and Relevance
Table 4. GPU efficiency comparison of ranking approaches. Throughput is measured as maximum items scored per GPU per second under a 500 ms latency budget.
Model QPS (Items/s/GPU) Latency (ms) GPU
Full Text 290 <<500 H100
Summarized, Pruned 2,200 <<500 H100
MixLM 22,000 <<500 H100
Embedding Retrieval $1.6\times 10^{9}$ <<100 A100
Table 3 shows that embedding-only retrieval delivers the lowest ranking quality and is therefore suitable only as a coarse first-stage filter. Full-text LLM ranking achieves the highest NDCG@10 (0.9432) but is prohibitively expensive, scoring only 290 items/s/GPU. Summarized-text LLM ranking with pruning improves throughput to 2,200 items/s/GPU while sacrificing some quality (NDCG@10 = 0.9218).
[…]
@@ 6.2. Online A/B Test
Table 5. Online A/B Test Results
[…]
@@
Semantic Job Search (MixLM) +0.47%
Table 5 summarizes the findings. After deploying MixLM in LinkedIn’s AI Job Search and running extensive A/B tests, we observed: MixLM has on-par relevance metrics comparing with the existing limited-scale production baseline(summarized text LLM)  (Behdin et al., 2025). With the 10.0x gain in serving throughput under the same latency budget, MixLM empowered the full-traffic deployment of LLM-ranking-based Semantic Job Search and for the first time, consequently producing a significant 0.47% increase in Daily Active Users, compared with the traditional job search.
@@ 6.3. Ablation Analysis
To understand the contribution of individual components, we perform ablation studies on a smaller dataset that mirrors the distribution of the full training corpus.
@@ 6.3.1. Scaling Up
Table 6. Ablation on Dataset Size
[…]
@@
Table 7. Embedding Tokens per Item
[…]
@@
Increasing both training data size and embedding granularity consistently improves ranking quality. For production, we retain a single embedding token per item to meet latency constraints.
@@ 6.3.2. Domain Reasoning Fine-Tuning
Table 8. Effect of Domain Reasoning Fine-Tuning
[…]
Vanilla LLM
@@
Domain reasoning fine-tuning substantially improves ranking quality and is adopted as the first stage of our training pipeline.
@@ 6.3.3. Self-Alignment and Auxiliary Losses
Table 9. Ablation on Auxiliary Losses
[…]
@@
+self-alignment +0.0014
+distillation +0.0091
+self-alignment + distillation +0.0108
Distillation provides the dominant gain, while self-alignment further improves results when combined with distillation.